Circular No. 7935

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dz IN MCG –01-1-52 B. Beutler and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on July 16.5 (mag ∼ 18.3) and 17.5 UT (mag ∼ 18.4). SN 2002dz is located at α = 0^{h}13^{m}34^{s}.08, δ = −5^{o}5′45′′.7 (equinox 2000.0), which is 4′′.6 west and 9′′.6 south of the nucleus of MCG –01-1-52. A KAIT image of the same field on 2001 Nov. 10.2 showed nothing at this position (limiting mag ∼ 19.0).

COMET 57P/DU TOIT-NEUJMIN-DELPORTE Y. R. Fernandez, D. C. Jewitt, and S. S. Sheppard, University of

Hawaii, report the detection of an additional 18 components (C–T) of comet 57P (cf. IAUC 7934) in observations taken on July 17.5 and 18.4 UT with the University of Hawaii 2.2-m reflector. R-band magnitudes range from 20.0 to 23.5. The components range from well condensed to diffuse with little central condensation and have comae of diameter 1′′–5′′. Components I, K, L, N, P, and S show a lack of central condensation. The components are delayed with respect to T = 2002 July 31.181 TT for component A by the following times (in days): C, +0.012; D, +0.037; E, +0.053; F, +0.078; G, +0.156; H, +0.164; I, +0.170; J, +0.180; (B, +0.188); K, +0.194; L, +0.194; M, +0.224; N, +0.226; O, +0.240; P, +0.271; Q, +0.309; R, +0.311; S, +0.313; and T, +0.354.

COMETS C/2002 M3, 2002 M4, 2002 M5, 2002 M6, 2002 M7 (SOHO)

Further to IAUC 7931, D. Hammer reports measurements for five additional Kreutz sungrazing comets, found on SOHO website C2 images by R. Kracht (C/2002 M3), X. Leprette (C/2002 M4), T. Hoffman (C/2002 M5), and R. Matson (C/2002 M6, C/2002 M7). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-N41.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 M3 June 23.746  6^{h}03.^{m}5  +21^{o}47′
C/2002 M4  27.064  6 17.3  +21 45
C/2002 M5  27.996  6 20.4  +21 45
C/2002 M6  28.854  6 23.8  +21 52
C/2002 M7  28.881  6 24.0  +21 55
2002 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green